OpenStudy (anonymous):

For f(x)= x^4+25x^2 + 144 answer the following questions. a) How many zeros will f(x) have? b) Find all the zeros of f(x). c) Factor f(x). d) What are the x intercepts of f(x)? USE A GRAPHING CALCULATOR TO CHECK YOUR ANSWERS

OpenStudy (anonymous):

If it says to use a graphing calculator to check your answers, maybe you should use a graphic calculator?

OpenStudy (phi):

the highest exponent is 4, so you will have 4 zeros you could find them by letting y= x^2 (so y^2 = x^4). sub into the original equation to get x^4+25x^2 + 144 =0 y^2 + 25y + 144 =0 solve for y using the quadratic formula then solve for x using \( x= \pm \sqrt{y} \) on both solutions for y.